OpenStudy (anonymous):

HELP ME PLEASE Jake wrote this equation for the area of a rectangle. 431.2 = 24.5w. What is the width (w) of this rectangle? A. 67.7 units B. 10,564.4 units C. 17.6 units D. 406.7 units

OpenStudy (agl202):

to find width, it would be very easy to have an expression defined by only w this can be done by either dividing both sides by 24.5 or subtracting 23.5w from both sides subtracting both sides by 23.5w will yield w = 431.2 -23.5w this doesn't really help so let's divide both sides by 24.5 24.5w/24.5 = 431.2/24.5 w = 17.6

OpenStudy (agl202):

Hope this helps!

OpenStudy (anonymous):

Thank you so much :) can you help me with a another one

OpenStudy (agl202):

New question plz...ur welcome!